OpenStudy (anonymous):

WILL GIVE MEDAL FOR ALGEBRA 2 HELP

OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (−2, 6) and has a slope of −1?

OpenStudy (anonymous):

use the formula y-y1=m(x-x1) where m is the slope and (x1,y1) is the given point. @bblock5

OpenStudy (anonymous):

i didnt get the right answer

OpenStudy (anonymous):

okay see here m=-1 and (x1,y1)=(-2,6) so the required answer is y-6=-1(x+2) or, y-6=-x-2 or x+y=4 @bblock5